Majorn Auto Parts Store had net income of $86,000 for the year just ended. Majorn collected the following additional information to prepare its statement of cash flows for the year:

Increase in accounts receivable $ 112,000
Decrease in inventory $ 23,000
Decrease in accounts payable $ 40,000
Increase in retained earnings $ 34,000
Cash received from sale of building $ 220,000
Gain on sale of building $ 52,000
Depreciation expense $ 37,000

Majorn uses the indirect method to prepare its statement of cash flows. What is Majorn's net cash provided by (used in) operating activities?


Multiple Choice

  • $46,000

  • $(58,000)

  • $200,000

  • $294,000
Answer #1

Answer:- $(58,000)

Statement of cash flows in indirect method-

Start with net income

>the non cash expenses and losses must be added back in and the gains must be subtracted.

>asset account increases-subtract the amount from the income

>asset account decreases-add amount to income

>liability account increases-add amount to income

>liability account decreases-subtract amount from income

so,  $86,000+37,000+23,000=146,000-112,000-52,000-40,000=$(58,000)